OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form. -x + 4y = -1 x - 4y = -1 x + y = 3 x - y = -3

OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form. Choices right here? -x + 4y = -1 x - 4y = -1 x + y = 3 x - y = -3

OpenStudy (anonymous):

ANYONE HERE EYYYYYYYYYYYYYYYYYY!

OpenStudy (anonymous):

EYYYYYYYYYYYYYYYYYYYYYYYYYYYYYY

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

Directrix (directrix):

(-1,4) is ( x1, y1) in the equation below Slope - Point Equation of a Line y - y1 = m ( x - x1) y - 4 = (-1)* ( x - (-1) ) y - 4 = (-1) * (x + 1) y - 4 = -x - 1 x + y + 1 - 4 = 0 x + y -3 = 0 And ..... Your Task --> Get the equation in standard form as defined above.

OpenStudy (anonymous):

I just did the math its x+y=3

Directrix (directrix):

Check out this online calculator of the equation of a line given a point on the line and the slope of the line: http://ncalculators.com/geometry/point-slope-form-calculator.htm

OpenStudy (anonymous):

so is this not the right answer?